4(2X - 3)= 2(X + 3)
Answer by Cromlix(4381)

4(2X - 3)= 2(X + 3)
Multiply out
8x - 12 = 2x + 6
collect x terms on left
numbers on right
Remember if a number or
x value changes side
it changes sign.
8x - 2x = 6 + 12
6x = 18
x = 3
Hope this helps
:-}